OpenStudy (ashleyy_smith):

At the city museum, child admission is $5.80 and adult admission is $9.40. On Friday, 143 tickets were sold for a total sales of 1013.00. How many child tickets were sold that day?

OpenStudy (anonymous):

child is c, costs 5.80 per ticket adult is a, costs 9.40 per ticket 5.80c+9.40a=1013 this finds the total amount of $ c+a=143 this finds the total amount of tickets sold you can use substitution to solve for children's tickets (c) a=143-c substitute (143-c) for a into the first equation. 5.80c+9.40(143-c)=1013 5.80c+1344.2-9.40c=1013 subtract 1344.2 from both sides combine the c terms -3.6c=-331.2 divide both sides by -3.6 c=92

OpenStudy (ashleyy_smith):

can u help me with 2 more? :o @ijakez

OpenStudy (anonymous):

go for it

OpenStudy (anonymous):

but so you can answer them without my help, you understand substitution right?

OpenStudy (ashleyy_smith):

yes :O

OpenStudy (ashleyy_smith):

Linda's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Linda $4.10 per pound, and type B coffee costs $5.45 per pound. This month's blend used four times as many pounds of type B coffee as type A, for a total cost of $725.20. How many pounds of type A coffee were used?

OpenStudy (anonymous):

coffee a can be defined as "a" coffee b can be defined as "b" what are the two equations about the relationship between coffee a and coffee b, and the relationship between both coffees and the cost?

OpenStudy (ashleyy_smith):

$4.10a and $5.45b

OpenStudy (anonymous):

good, but those together equal what? and what's the other equation?

OpenStudy (ashleyy_smith):

=725.20

OpenStudy (ashleyy_smith):

@ijakez

OpenStudy (anonymous):

alright, so 4.10a+5.45b=725.20 do you know the second equation or do you need help? it's the equation you determine from the statement "This month's blend used four times as many pounds of type B coffee as type A".

OpenStudy (ashleyy_smith):

i need help :s

OpenStudy (ashleyy_smith):

@ijakez

OpenStudy (anonymous):

all good! so it says that four times as much pound B coffee was used as pound A those words converted to an equation would look like 4b=a now that you have the 2 equations, 4b=a and 4.10a+5.45b=725.20, solve for a!

OpenStudy (ashleyy_smith):

im getting 35.6363636 wtf.

OpenStudy (anonymous):

substitute a with 4b: 4.10(4b)+5.45b=725.2 multiply 4.1 and 4b: 16.4b+5.45b=725.2 add the b's: 21.85b=725.2 divide both sides by 21.85 to solve for b: b=33.1899314 plug it back in to the equation 4b=a 4(33.1899314)=a a=132.7597256 a bit more calculator work, you can round to 3 decimals

OpenStudy (ashleyy_smith):

so it's 132.76? :o @ijakez

OpenStudy (anonymous):

yeah!

OpenStudy (ashleyy_smith):

I got it wrong :s @ijakez

OpenStudy (anonymous):

hmm, let me check my work then

OpenStudy (anonymous):

oh crap, it's 4a=b my mistake! can you resubmit?

OpenStudy (ashleyy_smith):

yes

OpenStudy (ashleyy_smith):

but I think it's ganna tell me to do 3 more problems now :/ since I got it wrong

OpenStudy (anonymous):

I'll just solve them quickly, my bad what are they?

OpenStudy (ashleyy_smith):

A Web music store offers two versions of a popular song. The size of the standard version is 2.2 megabytes (MB). The size of the high-quality version is 4.5 MB. Yesterday, the high-quality version was downloaded three times as often as the standard version. The total size downloaded for the two versions was 3297 MB. How many downloads of the standard version were there?

OpenStudy (ashleyy_smith):

@ijakez

OpenStudy (anonymous):

standard: a high quality: b b=3a 2.2a+4.5b=3297 solve for a 2.2a+4.5(3a)=3297 2.2a+13.5a=3297 15.7a=3297 a=210

OpenStudy (ashleyy_smith):

At the city museum, child admission is $6.10 and adult admission is $9.30. On Sunday, 163 tickets were sold for a total sales of $1227.90. How many adult tickets were sold that day? @ijakez

OpenStudy (anonymous):

child is a adult is b a+b=163 6.1a+9.3b=1227.9 a=163-b 6.1(163-b)+9.3b=1227.9 994.3-6.1b+9.3b=1227.9 994.3+3.2b=1227.9 3.2b=233.6 b=73

OpenStudy (ashleyy_smith):

1 more

OpenStudy (ashleyy_smith):

Latoya's Coffee Shop makes a blend that is a mixture of two types of coffee. Type A coffee costs Latoya $4.10 per pound, and type B coffee costs $5.30 per pound. This month, Latoya made 137 pounds of the blend, for a total cost of $630.10. How many pounds of type A coffee did she use?

OpenStudy (ashleyy_smith):

@ijakez

OpenStudy (anonymous):

a is a b is b 4.1a+5.3b=630.1 a+b=137 find a b=137-a 4.1a+5.3(137-a)=630.1 4.1a+726.1-5.3a=630.1 -1.2a=-96 a=80 done! so sorry

OpenStudy (ashleyy_smith):

awww its ok :o Dont feel bad! atleast u finished it for me :)